OpenStudy (anonymous):

Solve for the unknown x/100=4.8/1.7

sam (.sam.):

x/100=4.8/1.7 multiply both sides by 100 x=480/1.7 multiply both numerator and denominator by 10 x=4800/17

OpenStudy (anonymous):

The answer comes out to 282.35?

OpenStudy (anonymous):

I does not go with the answer from the book which is 320?

OpenStudy (anonymous):

Sometimes the book CAN be wrong. Ive come across a few wrong answers in the book & the teacher told the class. Make sure its the right answer to the question, right section & right problem juuust to be sure. 2.8235 or 282.35 should be right

OpenStudy (anonymous):

Oh. That's fine, but thanks for helping.

OpenStudy (anonymous):

No problem. Sorry we couldnt help much

sam (.sam.):

x=4800/17 That's in fraction , if you change it to decimal you'll get 282.35

OpenStudy (anonymous):

ok, .Sam., thanks.